Determine the Henry's law constant for ammonia in water at 35°C if an ammonia pressure of...

  1. Determine the Henry's law constant for ammonia in water at 35°C if an ammonia pressure of 0.56 atm produces a solution with a concentration of 0.67 M.

    1.2 M/atm

    0.029 M/atm

    59 M/atm

    7.9 M/atm
Answer #1

As per Henry law, solubility of gas is ditedire proportional to its partial pressure,

Equation is S = solubility = Kh*P

Kh is henry's law constant, P = pressure

Kh = S/P = 0.67/0.56 = 1.196 ~ = 1.2 M/atm

Correct answer : first option